How long does it take to send a 8 MiB file from Host A to Host B over a circuit-switched network, assuming: Total link transmission rate = 99 Gbps. Network is TDM, with 15 permitted users, each with an equal time slot size. A link connection requires a setup time of 92.5 ms. Your answer should be in miliseconds (ms) with one decimal place, and without the unit (e.g. "140.6" without the quotes). Show steps to get to 102.7.
Answer #1

Each circuit has a transmission rate = 99Gbps / 15 = 6.6 Gbps.

File size = 8 MiB = 8*1024*1024 bytes = 8*1024*1024*8 bits

Therefore, time to transmit the file = (8*1024*1024*8)/(6.6*106) ms = 10.2 ms.

Circuit establishment time = 92.5 ms

Total time = 92.5 + 10.2 = 102.7 ms